# Probing Einstein-Maxwell-Scalar Black hole via Thin Accretion Disks and Shadows with EHT Observations of M87* and Sgr A*

### v2

@@ I Introduction
Further advancements in EHT observations led to the polarization imaging of M87*, revealing the presence of a magnetic field, which provides insights into jet formation mechanisms 131 ; 132 ; 133 . Additionally, EHT reported the detection of the shadow of SgrA*, the supermassive BH at the center of the Milky Way 134 . These observations offer valuable constraints on the geometry of spacetime in the strong-field regime, particularly near the event horizon, allowing precise measurements of BH mass and spin 135 ; 136 ; 145 ; 146 ; 147 ; 148 ; 149 ; 150 ; 151 ; 152 ; 153 ; 154 ; 155 ; 185 . Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 .
@@
Among various modified gravity models, the Einstein-Maxwell-Scalar (EMS) model serves as a promising candidate. Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest. Previous studies have shown that dilaton BH spacetimes are influenced by the dilaton charge, which not only serves as a new form of hair but also provides a powerful tool for analyzing optical phenomena in such backgrounds. In this work, we consider a charged BH solution within the EMS theory as an extension of the Reissner-Nordström (RN) solution. Unlike dilaton BHs, where the scalar field typically contributes through an effective charge, the BH solution in EMS theory is directly governed by its model parameters, which serve as the primary distinguishing feature. A comprehensive exploration of the physical and optical properties of this solution is therefore essential. Investigating the effects of its parameters on spacetime geometry and optical phenomena will contribute to a deeper understanding of its astrophysical implications and provide a means to differentiate it from other BH solutions.
@@
In this article, we analyzed the optical properties of this BH solution through its shadow and the observable appearance of accretion disk emission in the strong-field regime. We further investigated the impact of EMS BH parameters on optical phenomena and constrained the viable ranges of parameters $\alpha$ and $\beta$ using EHT observational data. Our paper is organized as follows. In Sec.II, we will give the BH solution in EMS theory. In Sec.III, we will discuss the influence of EMS parameters on the optical properties of BH and plot the photon orbits of BH. Furthermore, we investigated the observable effects of radiation emitted by the accretion disk and found that the EMS BH exhibits a significantly reduced radiation intensity compared to the Schwarzschild and RN BHs. Finally, we present our conclusions and provide an outlook in Sec. IV. The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@ II The black hole solution in Einstein-Maxwell-scalar theory
The interaction between a scalar field and gravity was first systematically studied by Fisher, who obtained a static and spherically symmetric solution to the Einstein equations in the presence of a massless scalar field 102 . This pioneering work laid the foundation for subsequent investigations into the role of scalar fields in gravitational theories. In particular, the EMS theory has garnered considerable interest due to its profound physical implications and its natural emergence in various theoretical frameworks. Notably, it arises in the context of Kaluza-Klein theories, where extra-dimensional reductions lead to scalar field interactions 103 . Furthermore, it plays a crucial role in supergravity and string theory, where dilaton fields naturally couple to gravity and gauge fields 104 . Additionally, scalar fields are fundamental in cosmology, influencing early universe dynamics, inflationary models, and the late-time acceleration of the universe 105 . These diverse connections underscore the significance of the EMS theory in modern theoretical physics.
Within the framework of effective field theory (EFT), the interaction between a scalar field and an electromagnetic field can be incorporated through specific coupling terms in the Lagrangian. Such couplings naturally arise in the low-energy effective description of string theory, where the scalar field is identified with the dilaton—a massless excitation resulting from the compactification of extra spatial dimensions—while the electromagnetic field originates from gauge fields in the higher-dimensional theory. The dilaton, as a key modulus field, encodes information about the compactification geometry and the associated moduli space, dynamically influencing the strength of gauge interactions in the four-dimensional effective theory.
The coupling between the scalar and electromagnetic fields is an inevitable consequence of dimensional reduction, where the dilaton acts as a dynamical field that rescales the effective gauge couplings. This interaction not only modifies the propagation and self-interactions of electromagnetic waves but also introduces additional curvature terms in the gravitational sector, leading to non-trivial corrections to the Einstein field equations when the scalar field is present. These effects are particularly relevant in scenarios involving strong-field gravity, early-universe cosmology, and potential deviations from classical electrodynamics in high-energy regimes. Further implications include modifications to BH solutions, cosmological inflation models, and the possible emergence of new observational signatures in astrophysical and laboratory settings 169 ; 170 ; 171 ; 172 ; 173 ; 174 ; 175 ; 186 .
Such interaction terms are often introduced to match experimental or astrophysical observations under specific conditions. A typical form of the interaction term is given by
$\mathcal{L}_{\text{int}}=f(\phi)F_{\mu\nu}F^{\mu\nu},$ (1)
where $f(\phi)$ is a function of the scalar field $\phi$, and $F_{\mu\nu}$ represents the electromagnetic field strength tensor. Depending on the functional form of $f(\phi)$, the coupling may be constant, exponential, or take other more complex forms. These interaction terms can significantly influence the propagation of electromagnetic fields, especially in regions with strong scalar field backgrounds, such as near BHs or on cosmological scales.
Furthermore, the inclusion of a scalar field can be interpreted as a modification to standard Einstein gravity. Such frameworks are often categorized as scalar-tensor theories or extended theories of gravity. In these models, the coupling between the scalar field and the electromagnetic field can significantly influence the dynamics of the electromagnetic field and modify the solutions of the gravitational field equations. For example, investigations within the EMS theory have demonstrated that the presence of a scalar field can alter the electromagnetic properties of BHs, giving rise to novel classes of BH solutions, such as those characterized by scalar-electromagnetic coupling.
We consider the EMS theory described by the following action
$S[g_{\mu\nu},\phi]=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}(R-2\nabla_{\mu}\phi\nabla^{\mu}\phi-K(\phi)F_{\mu\nu}F^{\mu\nu}-V(\phi)),$ (2)
where $R$ is the Ricci scalar. The EMS theory describes the real scalar field $\phi$ minimally coupled to GR and non-minimally coupled to Maxwell’s background. $K(\phi)$ is the coupling function between the Maxwell and scalar field, and $V(\phi)$ is the scalar potential. if we set up the determined expressions of $K(\phi)$ and $V(\phi)$ first, then the corresponding BH solutions are determined. For example, if we set $K=1$ and $V=2\lambda$, $\lambda$ is the cosmological constant, then the EMS theory gives the Reissner-Nordstrom-de-Sitter solution. And if we set $K=e^{2\phi}$ and $V=0$, then we can get the dilation BH solution 1 ; 2 . Some important solutions with different $K(\phi)$ and $V(\phi)$ are given by 3 ; 4 ; 5 ; 6 ; 7 ; 8 .
In our research, we consider $V=0$ and the coupling function $K(\phi)$ is
$K(\phi)=\frac{(\alpha^{2}+1)e^{\frac{-2\phi}{\alpha}}}{(\alpha^{2}+1+\beta)e^{\frac{-2\phi(\alpha^{2}+1)}{\alpha}}+\beta\alpha^{2}},$ (3)
with $\alpha$ and $\beta$ as characteristic parameters or coupling constants; it is observed that when $\beta\to\infty$, the influence of the electromagnetic field is negligible. Conversely, in the extremal case where $\beta=0$, it simplifies to the Einstein-Maxwell-Dilaton (EMD) coupling described by $K(\phi)=e^{2\alpha\phi}$. The equations of motion are obtained by varying the action with respect to the metric, dilaton, and Maxwell field, respectively, as follows
@@
$\left\{\begin{array}[]{l}0=\nabla_{\mu}\left[K(\phi)F^{\mu\nu}\right],\\ \Box\phi=\frac{1}{4}\frac{\partial K(\phi)}{\partial\phi}F^{2},\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+2K(\phi)\left(F_{\mu\sigma}F^{~{}\sigma}_{\nu}-\frac{1}{4}g_{\mu\nu}F^{2}\right).\end{array}\right.$ (4)
Due to the field equations mentioned above, Ref.9 provides a BH solution, which can be described as
@@
$\left\{\begin{split}&ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)(d\theta^{2}+\sin^{2}\theta d\phi^{2})\\ &f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)}\\ &C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{split}\right.$ (5)
where $b_{1}$ and $b_{2}$ are functions solely of $\alpha$, yielding
$\left\{\begin{split}b_{1}&=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}&=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{split}\right.$ (6)
with $q\equiv{Q}/{M}$ the charge-to-mass ratio, and $M$ the mass of the BH.
The location of horizons is defined where $f(r_{\pm})=0$. Specifically, for $\beta=0$, these configurations align with the GMGHS solutions, as referenced in 11 , where $r_{\pm}$ are identified as $b_{1}$ and $b_{2}$, respectively. Additionally, when $\alpha=0$, the solution represent the horizon of RN BH. Conversely, setting $q=0$ simplifies the solution of the Schwarzschild BH. Notably, in EMS model, the horizons emerge strictly under the conditions $0<q<\sqrt{2}$ and $0<\beta<\frac{(2-q^{2})^{2}}{4q^{2}}$ 10 ; 45 .
@@
Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$,$\alpha$ and $q$.
@@
From the above two pictures, we can see there are two intersection points with $y=0$ which means the existence of the inner and outer horizons of BH. However, we see the third picture (EMD), when $\alpha>1$, there is only one intersection point with $y=0$. When the parameter $\alpha$ lies within the interval $(0,1)$, the EMD black hole possesses two horizons. Furthermore, we find that for a fixed $q$, an increase in $\alpha$ causes the function’s minimum value to decrease, whereas an increase in $\beta$ raises the minimum value. This may be attributed to the specific structure of the metric.
@@ III Null geodesics and the Shadow
In this section, we will first explore the behavior of the shadow radius theoretically, and then find the constraints to the EMS parameter using the data from EHT. To begin with, consider a static, spherically symmetric spacetime given by
$ds^{2}$ $=g_{\mu\nu}dx^{\mu}dx^{\nu}$ (7)
$=-A(r)dt^{2}+B(r)dr^{2}+C(r)d\Omega^{2}$
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the line element of the unit two-spheres. Without loss of generality, we analyze the null geodesic in the equatorial plane only such that the polar angle is fixed to $\theta=\pi/2$. Then, the Hamiltonian for light ray is given by
@@
It is worth noting that since we consider EMS BH, the metric functions satisfy $A(r)=f(r)$, $B(r)=f(r)^{-1}$, and $C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$. The equations of motion for null particles are
$\left\{\begin{split}&\dot{x}^{i}=\frac{\partial H}{\partial p_{i}},\\ &\dot{p}_{i}=-\frac{\partial H}{\partial x^{i}}.\\ \end{split}\right.$ (9)
Here, $\dot{x}=dx/d\lambda$ and $\dot{p}$ represents the conjugate momenta. Eq.(9) gives
$\left\{\begin{split}&\dot{t}=-\frac{p_{t}}{f(r)},\quad\dot{\phi}=\frac{p_{\phi}}{C(r)},\quad\dot{r}=-\frac{p_{r}}{f(r)^{-1}},\\ &\dot{p}_{t}=0,\quad\dot{p}_{\phi}=0,\\ &\dot{p}_{r}=\frac{1}{2}\left(-\frac{p_{t}^{2}f^{\prime}(r)}{f(r)^{2}}+\frac{p_{r}^{2}f^{\prime}(r)^{-1}}{f(r)^{-2}}-\tilde{C}p_{\phi}^{2}\right).\\ \end{split}\right.$ (10)
@@
$-\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (11)
@@
Setting $p_{t}=-\omega_{0}$, and using $p_{r}$, we get the relation how $r$ changes with $\phi$:
@@
For a circular light orbit, the radial velocity and acceleration should be $\dot{r}=0$ and $\ddot{r}=0$ respectively, and hence, $p_{r}=0$. Eq.(11) then becomes
@@
Since $\dot{p}_{r}=0$, Eq.(14) can be rewritten as
$\dot{p}_{r}=0=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}.$ (15)
Using Eq.(14) and (15), we find
$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)},$ (16)
and
$p^{2}_{\phi}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}}.$ (17)
The implication of subtracting Eqs. (16) and (17) give the information on how to find the radius of the photon sphere 176 ; 184
@@
Let us now determine the behavior of the shadow radius. For constructing the shadow, we assume that a static observer at radius coordinate $r_{\text{o}}$ sends light rays into the past. As can be seen from Fig.3, the angle $a$ between such a light ray and the radial direction is given by
$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}$ (19)
With the help of Eq.(12), we obtain
$\cot^{2}a=\frac{h\left(r_{\mathrm{O}}\right)^{2}}{h(R)^{2}}-1.$ (20)
By elementary trigonometry, we get 176
$\sin^{2}a=\frac{h(R)^{2}}{h\left(r_{\mathrm{O}}\right)^{2}},$ (21)
where $R$ means the minimum radius that the light ray approaches the center and then goes out.
The boundary curve of the shadow corresponds to past-oriented light rays that asymptotically approach one of the unstable circular light orbits at radius $r_{\mathrm{ph}}$. Therefore we have to consider the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(21) for getting the angular radius $a_{\text{sh }}$ of the shadow 184 ,
@@
Therefore we can also write Eq.(22) as
@@
For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{o}$ is sufficiently large, the black hole shadow radius approaches $b_{c}$. This radius is determined by the location of the photon sphere and the specific form of the metric function $f(r)$. In our paper, we will obey this consideration.
Figure 3: Example of calculation of light ray emitted from the observer’s position (small disk) into the past under an angle $a$. The BH horizon and the photon sphere are shown; $r_{\mathrm{ph}}$ is the photon sphere radius. The trajectory is calculated in the Asymptotically flat spacetime, $R$ denotes the radius coordinate at the point of closest approach. The picture is indicated in Ref.176 .

### v3

@@ I Introduction
Further advancements in EHT observations led to the polarization imaging of M87*, revealing the presence of a magnetic field, which provides insights into jet formation mechanisms 131 ; 132 ; 133 . Additionally, EHT reported the detection of the shadow of SgrA*, the supermassive BH at the center of the Milky Way 134 . These observations offer valuable constraints on the geometry of spacetime in the strong-field regime, particularly near the event horizon, allowing precise measurements of BH mass and spin 135 ; 136 ; 145 ; 146 ; 147 ; 148 ; 149 ; 150 ; 151 ; 152 ; 153 ; 154 ; 155 ; 185 . Collectively, these findings provide a crucial platform for testing GR and exploring potential modifications to gravitational theories in extreme astrophysical environments 137 ; Liu:2024brf ; Yin:2025coq ; He:2022lrc ; Wu:2024sng ; Feng:2024iqj ; Feng:2023iha ; Feng:2022bst .
@@
Among various modified gravity models, the Einstein-Maxwell-Scalar (EMS) model serves as a promising candidate. In fact, various specific realizations of EMS models naturally arise as consistent truncations or low-energy limits of string theory. For instance, in heterotic string theory, compactification from ten to four dimensions in the Einstein frame yields an effective action containing the metric, a dilaton-like scalar field, and Abelian gauge fields with an exponential coupling of the form $f(\phi)F_{\mu\nu}F^{\mu\nu}$ — as in the well-known Gibbons–Maeda–Garfinkle–Horowitz–Strominger (GMGHS) solution Gibbons:1988 ; Garfinkle:1990qj . Similar structures occur in Type II string compactifications, where moduli fields (Kähler, complex structure, axion–dilaton) couple to gauge fields descending from the NS–NS and R–R sectors Polchinski:1998 ; Becker:2006 . In extended supergravity, particularly $\mathcal{N}=4$ and $\mathcal{N}=8$ theories, the gauge kinetic matrix $f_{IJ}(\phi)$ depends explicitly on scalar fields Cremmer:1979 ; deWit:1982 . Truncation to a single $U(1)$ vector and one scalar degree of freedom yields precisely the EMS form. The $\mathcal{N}=2$ STU model Duff:1996 provides another example, in which selecting a single scalar–vector sector produces an EMS-type action.
The EMS action serves as a powerful unifying framework, encompassing two distinct and mutually exclusive branches of black hole solutions: the Einstein–Maxwell–Dilaton (EMD) branch and the spontaneously scalarized Reissner–Nordström (RN) branch. The profound differences between these two scenarios are encoded solely in the local behavior of the coupling function $K(\phi)$ at $\phi=0$. The EMD limit, realized by choosing an exponential coupling $K(\phi)=\mathrm{e}^{2\alpha\phi}$, is characterized by $K(0)=1$ and $K^{\prime}(0)\neq 0$. This condition renders the electro-vacuum RN solution, where $\phi\equiv 0$, an invalid background. Consequently, EMD black holes are born with an intrinsic, non-trivial scalar field, often referred to as “dilaton hair” and they never bifurcate from a hairless RN seed. Their asymptotic behavior is distinct, approaching a dilaton asymptotic at infinity, and they can violate the RN extremal bound on the charge-to-mass ratio Garfinkle:1990qj ; belkhadria_2025 .
In contrast, the spontaneously scalarized RN branch emerges from a different class of coupling functions satisfying $K(0)=1$, $K^{\prime}(0)=0$, and $K^{\prime\prime}(0)<0$. This specific set of conditions permits the RN metric with $\phi\equiv 0$ as a legitimate and stable background in the absence of a strong electric field. However, in the presence of a sufficiently strong electric charge, a tachyonic instability can arise. The effective mass squared of the scalar field, $\mu_{\text{eff}}^{2}(r)=K^{\prime\prime}(0)\,Q^{2}/r^{4}$, becomes negative within a certain radial window. This instability triggers a second-order phase transition, where the hairless RN black hole spontaneously acquires scalar hair, bifurcating from the original RN branch at a critical charge Qiu_2021 . Asymptotically, these scalarized black holes return to the RN solution, as the scalar hair vanishes at infinity.
@@
Since the EMS theory allows for hairy BH solutions 143 ; 144 ; 145 ; 189 , where BHs can possess scalar hair, investigating their properties in the strong-field regime is of significant interest. Previous studies have shown that dilaton BH spacetimes are influenced by the dilaton charge, which not only serves as a new form of hair but also provides a powerful tool for analyzing optical phenomena in such backgrounds. In this work, we consider a charged BH solution within the EMS theory as an extension of the RN solution. Unlike dilaton BHs, where the scalar field typically contributes through an effective charge, the BH solution in EMS theory is directly governed by its model parameters, which serve as the primary distinguishing feature. A comprehensive exploration of the physical and optical properties of this solution is therefore essential. Investigating the effects of its parameters on spacetime geometry and optical phenomena will contribute to a deeper understanding of its astrophysical implications and provide a means to differentiate it from other BH solutions.
@@
In this article, we analyzed the optical properties of this BH solution through its shadow and the observable appearance of accretion disk emission in the strong-field regime. We further investigated the impact of EMS BH parameters on optical phenomena and constrained the viable ranges of parameters $\alpha$ and $\beta$ using EHT observational data. Our paper is organized as follows. In Sec.II, we will give the BH solution in EMS theory. In Sec.III, we will discuss the influence of EMS parameters on the optical properties of BH and plot the photon orbits of BH. Furthermore, we investigated the observable effects of radiation emitted by the accretion disk and found that the EMS BH exhibits a significantly reduced radiation intensity compared to the Schwarzschild and RN BHs. And in Sec III.E, we also explore the deflection angle in the weak gravitational lensing. Finally, we present our conclusions and provide an outlook in Sec.IV. The paper has been written with the unit system $G=M=c=1$, and metric signature (,+,+,+)(-,+,+,+).
@@ II The black hole solution in Einstein-Maxwell-scalar theory
The systematic investigation of scalar-gravity coupling was introduced by Fisher, who obtained static spherically symmetric solutions to Einstein’s equations with a massless scalar field 102 . This foundational work established the theoretical basis for subsequent developments, particularly in Einstein-Maxwell-scalar (EMS) theory which has attracted significant attention due to its implications and natural emergence in: (i) Kaluza-Klein dimensional reduction 103 , (ii) dilaton couplings in string theory 104 , and (iii) cosmological inflation and late-time acceleration 105 .
The EMS action takes the form
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left[R-2(\nabla\phi)^{2}-K(\phi)F^{2}-V(\phi)\right],$ (1)
where $R$ is the Ricci scalar, $\phi$ is the scalar field. $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ represents the electro-magnetic field tensor. The field equations derived from this action yield
@@
$\begin{cases}\nabla_{\mu}[K(\phi)F^{\mu\nu}]=0,\\ \Box\phi=\frac{1}{4}\left[\partial_{\phi}K(\phi)F^{2}+\partial_{\phi}V(\phi)\right],\\ R_{\mu\nu}=2\partial_{\mu}\phi\partial_{\nu}\phi+\frac{1}{2}g_{\mu\nu}V(\phi)+2K(\phi)(F_{\mu\sigma}F^{\sigma}_{~\nu}-\frac{1}{4}g_{\mu\nu}F^{2}).\end{cases}$ (2)
The coupling function $K(\phi)$ and potential $V(\phi)$ determine the theory’s physical content. Some important solutions with different $K(\phi)$ and $V(\phi)$ are given by 3 ; 4 ; 5 ; 6 ; 7 ; 8 . Notable special solutions include
  • RN-de Sitter BH $K(\phi)=1$, $V=2\Lambda$, $\phi=0$,
  • Dilaton BH $K(\phi)=e^{2\phi}$, $V=0$ 1 ; 2 .
Different choices of the coupling function $K(\phi)$ lead to qualitatively distinct physics. Polynomial forms (such as $K(\phi)=1+\alpha\phi^{2}$) preserve the RN branch but allow for spontaneous scalarization above a critical threshold, where the scalar field effectively screens the electromagnetic charge and produces multiple scalarized branches. Exponential couplings (such as $K(\phi)=e^{2\alpha\phi}$) correspond to dilatonic models in which RN solutions no longer exist and charged black holes necessarily carry scalar hair, with modified extremality and thermodynamics. Other bounded or interpolating forms regulate the growth of scalar hair or bridge between EMS-type and dilatonic behavior.
Our investigation focuses on the parameterized coupling function Yu_2021
$K(\phi)=\frac{(\alpha^{2}+1)e^{-2\phi/\alpha}}{(\alpha^{2}+1+\beta)e^{-2\phi(\alpha^{2}+1)/\alpha}+\beta\alpha^{2}},$ (3)
which exhibits two physically significant limits. (i) $\beta\to\infty$: Decoupled electromagnetic sector. (ii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$. (iii) $\beta\to 0$: Einstein-Maxwell-Dilaton (EMD) theory with $K(\phi)=e^{2\alpha\phi}$.
The physical meaning of $\alpha$ is that it measures the strength of the coupling between the dilaton scalar field and the Maxwell field. It first emerges in the EMD limit ($\beta=0$) and is inherited from low-energy string theory, where the dilaton couples universally to gauge fields. Parameter $\beta$ controls the relative contribution of the RN and dilaton branches in the metric. Regarding its origin, it is introduced to combine the RN and dilaton line elements into a single interpolating family, thereby extending the EMD class.
As shown in Ref.9 , these admit BH solution with metric components
@@
$\begin{cases}ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)d\Omega^{2},\\ f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (4)
where the parameters $b_{1}$ and $b_{2}$ are determined by
$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M,\end{cases}$ (5)
with $q\equiv Q/M$ representing the dimensionless charge-to-mass ratio. Moreover, since obtaining analytic expressions for the inner and outer horizons of the EMS black hole is extremely difficult, we consider two special cases, while the general EMS case is determined numerically. (I) $\alpha=1$: With this definition established, the following relations can be derived Yu_2021
$\begin{cases}r_{+}=M\sqrt{2-q^{2}-\beta q^{2}+\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}},\\ r_{-}=M\sqrt{2-q^{2}-\beta q^{2}-\sqrt{\left(2-q^{2}\right)^{2}-4q^{2}\beta}}.\end{cases}$ (6)
For analytical convenience, we take $Q>0$ without imposing any loss of generality. The condition ensuring the existence of both the inner and outer horizons can be demonstrated to take the form
$0<q<\sqrt{2},\quad\quad\beta<\frac{\left(2-q^{2}\right)^{2}}{4q^{2}}.$ (7)
Moreover, the condition for the existence of a single horizon can be written as
$0<q<\sqrt{2},\quad\quad\beta=\frac{(2-q^{2})^{2}}{4q^{2}},$ (8)
under which the two horizons coincide, giving rise to an extremal black hole. This demonstrates that an extremal configuration can occur once the coupling constant $\beta$ meets the above criterion, irrespective of how small the charge is. (2) $\beta=0$ (the EMD solution): In this case, the inner and outer horizons are determined by $b_{1}$ and $b_{2}$.
[…]
The horizons are determined by the black hole’s mass $M$, charge $Q$, and the dilaton coupling constant $\alpha$. For two distinct real horizons to exist, the condition $q^{2}\leq 1+\alpha^{2}$ must hold. The equality $q^{2}=1+\alpha^{2}$ corresponds to the extremal case, where the two horizons merge into a single degenerate horizon. In contrast, the condition $q^{2}>1+\alpha^{2}$ eliminates the horizons and exposes a naked singularity. Therefore, when $\alpha\neq 1$ with arbitrary $\beta$, or when $\beta=0$ with arbitrary $\alpha$, a case-by-case analysis is necessary to ensure that naked singularities do not emerge in the black hole solutions. Notably, both situations correspond to special subclasses of the EMS framework. In the following analysis, we confine the parameters to these well-defined ranges to guarantee the existence of black hole solution.
Besides these, the black-hole horizons for all other parameter choices are already plotted in Figure 1. In the figure below, we plot the variation of the metric components of the EMS solution with respect to the coordinate radius $r$.
@@
Figure 1: The above three pictures plot the function $f(r)$ for different values of $\beta$$\alpha$ and $q$.
@@
From the above two pictures, we can see there are two intersection points with $y=0$ which means the existence of the inner and outer horizons of BH. However, in the third case (EMD), we observe that both a single intersection and two intersections can occur. All such parameter choices lie within the admissible ranges discussed above, thereby ensuring the absence of naked singularities. Furthermore, we find that for a fixed $q$, an increase in $\alpha$ causes the function’s minimum value to decrease, whereas an increase in $\beta$ raises the minimum value. This may be attributed to the specific structure of the metric.
@@ III Null geodesics and the Shadow
In this section, we will constrain the parameter of $\alpha$ and $\beta$ by comparing the theoretically defined shadow radius with observational data from the EHT. We begin by considering a static, spherically symmetric spacetime as presented in Ref.177
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (10)
The functions $A(r)$, $B(r)$, and $C(r)$ satisfy the asymptotic flatness conditions, with $A(r)$ and $B(r)$ $\rightarrow 1$ at spatial infinity while $C(r)\rightarrow r^{2}$. For simplicity, we restrict our analysis to null geodesics in the equatorial plane by fixing the polar angle at $\theta=\pi/2$. Under this configuration, the photon’s Hamiltonian takes the form
@@
It is worth noting that since we consider EMS BH, the metric functions satisfy $A(r)=f(r)$, $B(r)=f(r)^{-1}$, and $C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$. Subsequently, the equations of motion for null particles are
[…]
Using Eq.(12), we can derive
[…]
@@
$\frac{p_{t}^{2}}{f(r)}+\frac{p_{r}^{2}}{f(r)^{-1}}+\frac{p_{\phi}^{2}}{C(r)}=0,$ (14)
@@
Substituting the expressions for $p_{r}$ and $p_{t}=-\omega_{0}$, we finally obtain
@@
For a circular photon orbit, the conditions $\dot{r}=0$ (vanishing radial velocity) and $\ddot{r}=0$ (zero radial acceleration) necessarily imply $p_{r}=0$, which reduces Eq.(14) to the following form
@@
Furthermore, Eq.(17) can be expressed as
$\dot{p}_{r}=\frac{\omega_{0}^{2}f^{\prime}(r)}{f(r)^{2}}+\tilde{C}p^{2}_{\phi}=0.$ (18)
Using Eq.(17) and (18), we obtain
$p^{2}_{\phi}=C(r)\frac{\omega^{2}_{0}}{f(r)}=-\frac{\omega_{0}^{2}f^{\prime}(r)}{\tilde{C}f(r)^{2}},$ (19)
Combining Eq.(18) and Eq.(19), we give analytic expression for the photon sphere radius of the BH 176 ; 184 (See more details in Appendix A).
@@
Subsequently, the definition of the shadow radius must be established. The shadow radius behavior can be investigated by considering light rays traced backward in time from a static observer located at radial coordinate $r_{O}$. As shown in Fig.3, the angle $a$ formed between such light rays and the radial direction is expressed as
$\cot a=\left.\frac{\sqrt{g_{rr}}}{\sqrt{g_{\phi\phi}}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}=\left.\frac{\sqrt{B(r)}}{\sqrt{C(r)}}\frac{dr}{d\phi}\right|_{r=r_{\mathrm{O}}}.$ (21)
With the help of Eq.(15), we obtain 176
[…]
where $R$ means the minimum radius that the light ray approaches the center and then goes out (turning point of the trajectory). The shadow boundary is determined by past-oriented light rays asymptotically approaching the unstable circular photon orbits at radius $r_{\mathrm{ph}}$. Consequently, the angular radius $a_{\text{sh}}$ of the shadow is obtained by taking the limit $R\rightarrow r_{\mathrm{ph}}$ in Eq.(22) 184 .
@@
Therefore, the Eq.(23) can also be expressed as
@@
For a general static spherically symmetric asymptotically flat spacetime, Ref.176 demonstrates that when the observer’s radial coordinate $r_{O}$ is sufficiently large, the BH shadow radius approaches $b_{c}$. This radius is determined by the location of the photon sphere and the specific form of the metric function $f(r)$. In our paper, we will obey this consideration.
Figure 3: Example of calculation of light ray emitted from the observer’s position into the past under an angle $a$. The BH horizon and the photon sphere are shown; $r_{\mathrm{ph}}$ is the photon sphere radius. The picture is indicated in Ref.176 .